# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ III-C Energy-based model
[…]
The parameters in an energy-based model in Eq.(4) can be estimated by maximum likelihood estimation (MLE). Specifically, if $\mathbf{x}^{t}\sim p_{data}(\mathbf{x})$ for $t=1,2,...,n$, the likelihood function can be expressed as,
$\small\mathcal{L}(\theta)=\frac{1}{n}\sum_{t=1}^{n}\text{log}p_{\theta}(\mathbf{x}^{t})=\mathbb{E}_{\mathbf{x}\sim p_{data}(\mathbf{x})}[\text{log}p_{\theta}(\mathbf{x})].$ (5)
Since the energy function $g_{\theta}(\mathbf{x})$ in Eq.(4) can use different network architectures, it can approximate any kind of data distribution. In addition, the energy function can be multimodal without resorting to mixture distributions and does not differentiate between discrete and continuous distributions. These advantages make the energy-based model more flexible than traditional methods like kernel density estimation (KDE) [37].
IV Methodology
By definition of Shapley value in Eq.(1), it can be seen that the most important step is to estimate the contribution function $v(S)$, which can be reformulated as the conditional expectation in Eq.(3). The conditional expectation can be empirically approximated by techniques like Monte Carlo integration provided the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}^{t})$ is known.
[…]
where $\mathbf{x}_{\bar{S}}^{(k)}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ for $k=1,2,...,K$, and $K$ is the number of Monte Carlo sampling. It is worth noting that estimation of conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be achieved by techniques like KDE. However, KDE is computationally challenging for high-dimensional inputs. More importantly, using KDE requires estimation of $2^{|D|-1}$ conditional densities as there are a total of $2^{|D|-1}$ combinations of $S$.
IV-A Network structure
To achieve an accurate estimation of conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, an energy-based model called EmSHAP is presented here, which is capable of estimating all conditional densities using a single model. It has the network structure as shown in Fig. 1.
Figure 1: Structure of GRU coupled energy-based model for estimating Shapley value
In Fig. 1, data batches of input features are fed into the energy network to estimate the energy function $\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$, which can be used to approximate the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ after divided by the partition function. Here, the energy function $g_{\theta}$ can be designed as an unnormalized MLP with skip connections [38]. The partition function, on the other hand, needs to be approximated by Monte Carlo integration from the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$. The proposal distribution can be estimated using an autoregressive approach [39], however, its estimation accuracy will be negatively affected by feature orderings [^1] Feature ordering refers to the arrangement of input features in a specific sequence when fed into a neural network.. Instead, a GRU network with a dynamic masking scheme is used to estimate the proposal distribution, which outputs parameters related to the estimated proposal distribution and a context vector $\gamma$. The context vector is used to connect the energy network and the GRU network. With the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ obtained, the contribution function can be estimated from Eq.(6), and Shapley value can be subsequently obtained from Eq.(1). This network structure is under the framework of energy-based model, which can universally approximate any probability density according to Ref. [40]. Hence the proposed model structure inherits the good approximation ability of EBM. Comparing to the original EBM, a GRU network is used to generate the proposal distribution and integrate importance sampling to effectively estimate the partition function. The GRU sequentially generates the distribution of latent variables based on the masked variables and ultimately combines these into the proposal distribution. By capturing the intrinsic information associated with the sequential masking of variables, the GRU network ensures that the generated proposal distribution is more accurate, thereby enabling the model to better approximate complex conditional probability densities. The details of the network structure will be explained in Subsection IV-B.
IV-B GRU coupled energy-based model
When using the EBM to approximate the conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, we are interested in estimating the conditional density of inputs in $\bar{S}$ conditioned on the remaining observed inputs (in $S$) for all possible $\bar{S}$. According to Eq.(4), $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be approximated using the EBM as follows.
$\small p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\frac{\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))}{Z_{{\bar{S}};\mathbf{x}_{S}}}.$ (7)
@@ IV-B GRU coupled energy-based model
The right hand side of Eq.(7) can be transformed to a probability once $Z_{\bar{S},\mathbf{x}_{S}}$ is known. However, it is always intractable to directly calculate $Z_{\bar{S},\mathbf{x}_{S}}$. One solution is to introduce a proposal conditional distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ that is close to $Z_{\bar{S},\mathbf{x}_{S}}$, then use Monte Carlo integration to sample from $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to approximate $Z_{\bar{S},\mathbf{x}_{S}}$.
[…]
@@
$\small q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\prod_{l=1}^{|{\bar{S}}|}q(x_{{\bar{S}_{l}}}|\mathbf{x}_{S\cup\bar{S}_{<l}}).$ (9)
Eq.(9) uses the chain rule of conditional probability to obtain the proposal distribution, so that estimation of the proposal distribution can be converted into the multiplication of a series of one-dimensional conditional densities $q(x_{{\bar{S}}_{l}}|\mathbf{x}_{S\cup\bar{S}_{<l}})$, which is easy and tractable.
@@
Figure 2: GRU network for proposal distribution estimation
In Fig. 2, the unmasked features $\mathbf{x}_{S}$ together with the $j$-th masked feature $x_{\bar{S}_{j}}$ are fed into the GRU cell to generate the $j$-th hidden state $\mathbf{h}_{j}$ based on the information of the $j-1$-th hidden state $\mathbf{h}_{j-1}$, which can be summarized as follows.
@@
It should be noted that a total of $|D|-1$ zeros are combined with $x_{\bar{S}_{j}}$ in Fig. 2 to align with the dimension of $\mathbf{x}_{S}$. The hidden state has a dimension of $2|D|+|\bm{\gamma}|$, with the first $|D|$ elements forming the mean vector of the auxiliary variables $\mathbf{z}_{j}$, the second $|D|$ elements forming the diagonal vector of $\mathbf{z}_{j}$’s covariance matrix, and the last $|\bm{\gamma}|$ elements forming the context vector $\bm{\gamma}_{j}$. The hidden state $\mathbf{h}_{j}$ can be denoted as follows.
@@
Here $\bm{\mu}_{\mathbf{z}_{j}}$ is the mean vector, $\bm{\sigma}_{\mathbf{z}_{j}}$ is the vector containing the diagonal elements of $\mathbf{z}_{j}$’s covariance matrix. The context vector $\bm{\gamma}_{j}$ is used to connect the GRU network with the energy network.
The auxiliary variables $\mathbf{z}_{j}$ has a Gaussian distribution as follows.
@@
Another notable issue here is that Gaussian distribution is used as the prior distribution of the auxiliary variables. Although it is not a perfect prior for complex data or model, it is widely used in importance sampling. According to Ref. [41], importance sampling can effectively approximate the partition function as long as the proposal distribution has similar shape or overlapping region with the real distribution, and Gaussian distribution well fits this condition.
With the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ determined, the loss function of the GRU network can be established according to the maximum likelihood criterion.
$\quad\mathcal{J}({\Theta};\mathbf{x})=-\mathcal{L}({\Theta};\mathbf{x})$ (14)
@@
Here ${\Theta}$ is the parameter set.
Using the GRU network structure in Fig. 2, the conditional densities for any combination of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}$ can be obtained so that a single model can estimate all the conditional densities and repetitive operation of conditional density estimations can be avoided.
In addition, the difference in the variance of the data sampled from the proposal conditional distribution and the true conditional distribution has an explicit expression. Specifically, the data sampled from the true conditional distribution has the following variance,
$\small\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}]-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]\right)^{2}.$ (15)
Meanwhile, the variance of the data sampled from the proposal conditional distribution is,
$\quad\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]$ (16)
$\quad-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]\right)^{2}$
Subtract Eq.(15) and Eq.(16), the following equation is obtained:
$\quad\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]-\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]$ (17)
As can be seen from Eq.(17), when the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ is close to the true distribution $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, the discrepancy in variance between the approximate partition function $Z_{{\bar{S}};\mathbf{x}_{S}}$ and the true partition function approaches zero.
IV-C Dynamic masking scheme
Calculation of Shapley values requires estimating the conditional expectations of contribution function under different feature combinations. This involves repeatedly evaluating the impact on the model output by including or excluding specific input features. The masking scheme provides an effective way to perform this evaluating procedure. In the masking scheme, a Bernoulli distribution is used to determine whether an input feature is included or excluded, which can be described as follows.
@@ IV-C Dynamic masking scheme
Here, $\zeta$ is the masking probability, and $b_{i}$ is a binary value generated from the Bernoulli distribution, with $b_{i}=0$ indicating $x_{i}$ being included in the feature set $S$, otherwise $x_{i}$ is masked and excluded from the feature set $S$. The masking probability $\zeta$ is an important parameter that needs to be manually tuned, for example, $\zeta$ can be set as 0.5 by trial and error. However, using a fixed masking probability is problematic, as it may not generalize to a wide range of potential feature combinations, especially when the number of possible feature combinations is very large. Here, a dynamic masking scheme is introduced by gradually increasing $\zeta$, for example, from 0.2 to 0.8, to cover a wider range of feature combinations. Another benefit of using the dynamic masking scheme is that there is no need for manually tuning the optimal masking rate.
@@
By applying the dynamic masking scheme, the GRU network recursively processes the masked input features to capture dependencies among variables. At each step, GRU updates its hidden state by incorporating historical masking information and generates the conditional probability estimate for the current step. This process enables GRU to leverage its internal gating mechanisms to retain and propagate prior masking information, thus effectively capturing feature dependencies without relying on a fixed feature order. The detailed procedures for training EmSHAP are listed in Algorithm 1.
@@
        $\mathbf{h}_{j}=\text{GRU}\left(\left[\mathbf{x}_{S},x_{\bar{S}_{j}}\right],\mathbf{h}_{j-1}\right)$
        $\left[\bm{\mu}_{\mathbf{z}_{j}}^{T},\bm{\sigma}_{\mathbf{z}_{j}}^{T},\bm{\gamma}_{j}^{T}\right]^{T}=\mathbf{h}_{j}$
@@
     energy network$([\mathbf{x}_{\bar{S}}^{(\tilde{k})},\mathbf{x}_{S}],\bm{\gamma}_{|\bar{S}|})$rightarrow $\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$

### v4

@@ IV-B Energy network for estimation of unnormalized conditional probability density
where $Z=\int\exp(-g_{\theta}(\mathbf{x}))d\mathbf{x}$ is the partition function.
A key advantage of EBMs lies in their flexibility: by parameterizing the energy function $g_{\theta}(\mathbf{x})$ with neural networks, they can serve as a powerful framework for modeling complex probability distributions. In particular, this formulation allows EBMs to represent both joint and conditional distributions. Extending Eq.(5) to the conditional case yields,
$\small p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\frac{\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))}{Z_{\bar{S};\mathbf{x}_{S}}},$ (6)
where $\bar{S}$ denotes the unobserved subset and $S$ the observed subset of features. This conditional formulation is essential for Shapley value estimation, as it requires repeatedly modeling the conditional distribution of missing features given observed ones. Here, $Z_{\bar{S};\mathbf{x}_{S}}=\int\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))d\mathbf{x}_{\bar{S}}$ is the partition function specific to the subset $\bar{S}$ and observed features $\mathbf{x}_{S}$.
In our implementation, $g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S})$ is parameterized by a four-layer MLP with two residual connections [43], and can be regarded as an energy network. In the energy network, the skip connections help stabilize training and improve representation capacity, especially when modeling high-dimensional feature spaces. The input of the energy network consists of both the observed features $\mathbf{x}_{S}$ and the masked (unobserved) features $\mathbf{x}_{\bar{S}}$, allowing the network to capture interactions between observed and masked features. The energy network for estimating the unnormalized conditional probability density is as follows.
$\small g_{\theta}(\mathbf{x}_{S},\mathbf{x}_{\bar{S}})=\text{MLP}([\mathbf{x}_{S},\mathbf{x}_{\bar{S}}]).$ (7)
The energy network outputs a scalar energy value, where a lower energy corresponds to a higher probability. This scalar output is used to calculate the unnormalized probability $\exp(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$. After dividing this unnormalized probability by the partition function $Z_{\bar{S};\mathbf{x}_{S}}$, the normalized conditional probability of the unobserved features given the observed ones can be obtained.
IV-C GRU network for estimating the proposal distribution
@@
When using the EBM to approximate the conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, it is always intractable to directly calculate $Z_{\bar{S},\mathbf{x}_{S}}$. One solution is to introduce a proposal conditional distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ that is close to $Z_{\bar{S},\mathbf{x}_{S}}$, then use Monte Carlo integration to sample from $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to approximate $Z_{\bar{S},\mathbf{x}_{S}}$.
$\hat{Z}_{\bar{S};\mathbf{x}_{S}}$ $=\int\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))d\mathbf{x}_{\bar{S}}$ (8)
$=\int\frac{\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})d\mathbf{x}_{\bar{S}}$
@@
$\small q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})\approx\prod_{j=1}^{|{\bar{S}}|}q(x_{{\bar{S}_{j}}}|\mathbf{x}_{S\cup\bar{S}_{<j}}),$ (9)
where $j=1,\dots,|{\bar{S}}|$ is the index of features in $|S|$. Eq.(9) uses the chain rule of conditional probability to obtain the proposal distribution, so that estimation of the proposal distribution can be converted into the multiplication of a series of one-dimensional conditional densities $q(x_{{\bar{S}}_{l}}|\mathbf{x}_{S\cup\bar{S}_{<l}})$, which is easy and tractable.
@@
Figure 2: GRU network for proposal distribution estimation.
In Fig. 2, the unmasked features $\mathbf{x}_{S}$ together with the $j$-th masked feature $x_{\bar{S}_{j}}$ are fed into the GRU cell to generate the $j$-th hidden state $\mathbf{h}_{j}$ based on the information of the j1j-1-th hidden state $\mathbf{h}_{j-1}$, which can be summarized as follows.
@@
It should be noted that a total of |D|1|D|-1 zeros are combined with $x_{\bar{S}_{j}}$ in Fig. 2 to align with the dimension of $\mathbf{x}_{S}$. The hidden state has a dimension of $2|D|+|\bm{\gamma}|$, with the first $|D|$ elements forming the mean vector of the auxiliary variables $\mathbf{z}_{j}$, the second $|D|$ elements forming the diagonal vector of $\mathbf{z}_{j}$’s covariance matrix, and the last $|\bm{\gamma}|$ elements forming the context vector $\bm{\gamma}_{j}$. The hidden state $\mathbf{h}_{j}$ can be denoted as follows.
@@
Here $\bm{\mu}_{\mathbf{z}_{j}}$ is the mean vector, $\bm{\sigma}_{\mathbf{z}_{j}}$ is the vector containing the diagonal elements of $\mathbf{z}_{j}$’s covariance matrix. The context vector $\bm{\gamma}_{j}$ is used to connect the GRU network with the energy network by transferring the contextual information, thereby facilitating the interaction between the two components and enhancing the overall performance of conditional probability estimation. It does not explicitly involves in feature attribution, but rather provides insight into how feature dependencies are captured and transferred within EmSHAP. Table E2 of Appendix E analyzed how the length of $\bm{\gamma}$ impacts the estimation accuracy.
The auxiliary variables $\mathbf{z}_{j}$ have a Gaussian distribution as follows.
@@
Another notable issue here is that Gaussian distribution is used as the prior distribution of the auxiliary variables. Although it is not a perfect prior for complex data or models, it is widely used in importance sampling. According to Ref. [49], importance sampling can effectively approximate the partition function as long as the proposal distribution has a similar shape or overlapping region with the real distribution, and Gaussian distribution well fits this condition.
Eq.(10)-Eq.(13) naturally lead to the GRU network in Figs. 1 and 2. By substituting the obtained proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ into Eq.(8), the approximate partition function $\hat{Z}_{\bar{S};\mathbf{x}_{S}}$ can be obtained. This partition function is then incorporated into Eq.(6) to derive the conditional probability density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ based on the EBM. Finally, substituting $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ into Eq.(4) yields an approximate expression for the contribution function $v(S)$. These derivations provide the foundation for estimating the desired quantities using EmSHAP. With the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ determined, the loss function of EmSHAP can be established according to the maximum likelihood criterion.
$\qquad\mathcal{J}({\Theta};\mathbf{x})=-\mathcal{L}({\Theta};\mathbf{x})$ (14)
@@
Here $\Theta$} is the parameter set. Using the GRU network structure in Fig. 2, the conditional densities for any combination of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}$ can be obtained so that a single model can estimate all the conditional densities and repetitive operation of conditional density estimations can be avoided.
IV-D Dynamic masking scheme
Calculation of Shapley values requires estimating the conditional expectations of the contribution function under different feature combinations. This involves repeatedly evaluating the impact on the model output by including or excluding specific input features. The masking scheme provides an effective way to perform this evaluating procedure. In the masking scheme, a Bernoulli distribution is used to determine whether an input feature is included or excluded, which can be described as follows.
@@ IV-D Dynamic masking scheme
Here, $\zeta$ is the masking probability, and $b_{i}$ is a binary value generated from the Bernoulli distribution, with $b_{i}=0$ indicating $x_{i}$ being included in the feature set $S$, otherwise, $x_{i}$ is masked and excluded from the feature set $S$. The masking probability $\zeta$ is an important parameter that needs to be manually tuned, for example, $\zeta$ can be set as 0.5 by trial and error. However, using a fixed masking probability is problematic, as it may not generalize to a wide range of potential feature combinations, especially when the number of possible feature combinations is very large. Here, a dynamic masking scheme is introduced by gradually increasing $\zeta$, for example, from 0.2 to 0.8, to cover a wider range of feature combinations. Another benefit of using the dynamic masking scheme is that there is no need for manually tuning the optimal masking rate.
@@
   $\mathbf{h}_{j}=\text{GRU}([\mathbf{x}_{S},x_{\bar{S}_{j}}],\mathbf{h}_{j-1})$
   $\bm{\mu}_{\mathbf{z}_{j}}^{T},\bm{\sigma}_{\mathbf{z}_{j}}^{T},\bm{\gamma}_{j}^{T}]^{T}=\mathbf{h}_{j}$
@@
  energy network $g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S})=\text{MLP}([\mathbf{x}_{\bar{S}}^{(\tilde{k})},\mathbf{x}_{S}],\bm{\gamma}_{|\bar{S}|})$
@@
By applying the dynamic masking scheme, the GRU network recursively processes the masked input features to capture dependencies among variables. At each step, GRU updates its hidden state by incorporating historical masking information and generates the conditional probability estimate for the current step. This process enables GRU to leverage its internal gating mechanisms to retain and propagate prior masking information, thus effectively capturing feature dependencies without relying on a fixed feature order. The detailed procedures for training EmSHAP are listed in Algorithm 1, and the pipeline of using EmSHAP to calculate Shapley values is illustrated in Fig. 3.
Figure 3: Pipeline of EmSHAP for estimation of Shapley values.
IV-E Computational complexity analysis of EmSHAP
This subsection analyzes the computational complexity of EmSHAP. The computational burden of EmSHAP primarily arises from two components: the energy network and the GRU network. The energy network is implemented as a fully connected network with two residual connections, with a complexity of $O((2|D|+|\bm{\gamma}|)\cdot H+H^{2})$, where $H$ is the number of latent dimensions. The GRU network consists of a single GRU layer, whose complexity is $O(L\cdot(2|D|+|\bm{\gamma}|)^{2}+L\cdot(2|D|+|\bm{\gamma}|)\cdot 2|D|)$, where $L$ is the sequence length. Hence, the overall computational complexity of EmSHAP can be approximated as $O(L\cdot(2|D|+|\bm{\gamma}|)^{2}+L\cdot(2|D|+|\bm{\gamma}|)\cdot(2|D|+H)+H^{2})$. Since EmSHAP evaluates samples individually, the sequence length $L$ is typically set to 1, which simplifies the computation complexity to $O((2|D|+|\bm{\gamma}|)^{2}+(2|D|+|\bm{\gamma}|)\cdot(2|D|+H)+H^{2})$. It can be seen that the computational complexity is determined by the input feature dimension $|D|$, the latent dimension $H$, and context vector length $|\bm{\gamma}|$.